# Inflationary Flows: Calibrated Bayesian Inference with Diffusion-Based Models

### v1

@@ 7 Discussion
Here, we have proposed a new type of implicit probabilistic model based on the probability flow ODE (pfODE) in which it is possible to perform calibrated, identifiable Bayesian inference on a reduced-dimension latent space via sampling and integration. To do so, we have leveraged a correspondence between pfODEs and diffusion-based models by means of their associated Fokker-Planck equations, and we have demonstrated that such models continue to produce high-quality generated samples even when latent spaces are as little as 2% of the nominal data dimension. More importantly, the uniqueness and controllable error of the generative process make these models an attractive approach in cases where accurate uncertainty estimates are required.
Limitations: One limitation of our model is its reliance on the participation ratio (7) as a measure of dimensionality. Because PR relies only on second-order statistics and our proposals (9) are formulated in the data eigenbasis, our method tends to favor the top principal components of the data when reducing dimension. However, as noted above, this is not simply via truncation of lower PCs, since dimensions still mix via coupling to the score function in (6). Nonetheless, solutions to the condition (8) that preserve (or reduce) more complex dimensionality measures might lead to even stronger compressions for curved manifolds (Appendix C.2.1), and more sophisticated choices for noise and rescaling schedules in (6) might lead to compressions that do not simply remove information along fixed axes, more similar to [47]. That is, we believe much more interesting classes of flows are possible. A second limitation is that mentioned in Section and our experiments: our schedule requires training DBMs over much larger ranges of noise than are typically used, and this results in noticeable tradeoffs in compression performance as the number of dimensions is varied.
Related work: This work draws on several related lines of research, including work on using DBMs as likelihood estimation machines [29, 48, 22], relations with normalizing flows and hierarchical VAEs [48, 24, 49], and generative flow networks [50]. By contrast, our focus is on the use of DBMs to learn score functions estimates for implicit probabilistic models, with the ultimate goal of performing accurate inference. In this way, it is also closely related to work on denoising models [36, 37, 47, 51] that cast that process in terms of statistical inference and to models that use DBMs for de-blurring and in-painting [52, 53]. However, this work is distinct from several models that use reversal of deterministic transforms to train generative models [54, 55, 56, 57]. Whereas those models work by removing information from each sample $\mathbf{x}$, our proposal relies critically on adjusting the local density of samples with respect to one another, moving the marginal distribution toward a Gaussian.
@@ B.1.4 Re-writing loss with optimal preconditioning factors
To construct $\mathbf{g}$ in our PR-Reducing experiments, we chose to make all elements of $\mathbf{g}$ corresponding to preserved dimensions equal to 2 (i.e., $g_{preserved\_dims}=g_{max}=2$) and all elements corresponding to compressed dimensions equal to $g_{compressed\_dims}=g_{min}=1-\frac{d}{(D-d)}$, where $d$ is the number of dimensions being preserved and $D$ is total number of dimensions. Note that, using this construction, PR-Reducing schedules with a higher number of preserved dimensions $d$ have a smaller value of $\mathbf{g}$ for compressed dimensions, for a given fixed $D$. As such, PR-Preserving networks with higher $d$ are exposed to a higher noise range ratio (i.e., higher $e^{\rho(g_{max}-g_{min})t}$) during training and need to optimally amortize score estimation over this wider “inflation gap,” which is a harder learning problem. This is confirmed empirically by looking at the results of our Frechet Inception Distance (FID) [46] and round-trip mean squared error (MSE) experiments (Table 1), which show that PR-Reducing networks with higher number of preserved dimensions and, therefore, wider inflation gaps, have higher FID scores and MSEs.
We present values for $g_{preserved\_dims}$ and $g_{compressed\_dims}$ constructed using the prescription above for our toy and image experiments in Appendix B.4. Additionally, in Tables 3 and 7, we show the latent space variances achieved at the end of inflation for compressed dimensions across the different PR-Reducing regimes investigated here (i.e., PR-Reducing to 1D or 2D, for toys, and PR-Reducing (PRR) to 2%, 10%, and 20% of nominal dimensions, for images). Note that, as expected, the schedules with wider “exponential inflation gap” (i.e., PR-Reducing to 2D (toys) or to 20% (images)) achieve the highest effective compression (lowest latent space variance for compressed dimensions). Therefore, for our specific $\mathbf{g}$ construction scheme, actual compression achieved should be understood both in terms of number of dimensions preserved and in terms of effective latent space variance (for compressed dimensions) achieved at the end of inflation.
@@ B.4.1 Toy DataSets
Toy models were trained using a smaller convolutional UNet architecture (ToyConvUNet) and our proposed preconditioning factors (Appendix B.1). For all toy datasets, we trained networks both by using original images as inputs (i.e., “image space basis”) or by first transforming images to their PCA representation (i.e., “eigenbasis”). Networks trained using either base choice were able to produce qualitatively good generated samples, across all datasets. For all cases, we used $\rho=1$, a learning rate of $10^{-5}$, batch size of 8192, and exponential moving average half-life of $50\times 10^{4}$. Additionally, for PR-Reducing schedules, we chose $g_{preserved\_dims}=2$ (all datasets), and $g_{compressed\_dims}=0$ (2D datasets), or $g_{compressed\_dims}=-1$ (3D datasets) (cf. Appendix B.2). The only exceptions were networks used on mesh/ 2D alpha-shape toy experiments, where we used instead $g_{preserved\_dims}=1.15$ across all preserved dimensions (circles, S-curve) and $g_{compressed\_dims}=0.85$ (circles), or $g_{compressed\_dims}=0.70$ (S-curve). This yields a softer effective compression (i.e., smaller “inflation gap”) and is needed to avoid numerical underflow errors when fitting meshes or 2D alpha-shapes to the compressed latent spaces (see Appendix B.6).
As explained in Appendix B.1, to construct our $\mathbf{c}_{noise}$ preconditioning factor, we sampled $t\sim\mathcal{U}(t_{min},t_{max})$, with $t_{min}=10^{-7}$ across all simulations and $t_{max}$ equal to the values shown in Table 2. In the same table, we also show training times (in millions of images (Mimgs, as in [28])), along with both the total number of dimensions (in the original data) and the number of dimensions preserved (in latent space) for each dataset and schedule combination. In Table 3, we showcase latent space (i.e., end of “inflation”) compressed dimension variances achieved for the different toy PR-Reducing experiments.
@@
Table 3: Toy Experiments Latent Space Compressed Dimension Variance
Dataset Schedule Dimensions Kept Latent Space Compressed Dimension Variance
Circles PRR 1 $4\times 10^{-7}$
Sine PRR 1 $4\times 10^{-7}$
Moons PRR 1 $4\times 10^{-7}$
S Curve PRR 2 $2\times 10^{-12}$
Swirl PRR 2 $2\times 10^{-12}$
@@ B.4.2 CIFAR-10 and AFHQv2 Datasets
For our image datasets (i.e., CIFAR-10 and AFHQv2), we utilized similar training hyperparameters to the ones proposed by [28] for the CIFAR-10 dataset, across all schedules (e.g., PR-Preserving (PRP), PR-Reducing (PRR) to 2, 10, and 20%) (Table 4). Shown in Table 5 are our specific choices for the exponential inflation constant ($\rho$), for $g_{preserved\_dims}$ and $g_{compressed\_dims}$ values, training durations (in Mimgs), and both total and preserved number of dimensions per each schedule and dataset combination. Of note, for the PR-Preserving schedule, we tested two different configurations with smaller and larger noise. These differ by our choice of $\rho$, with smaller noise using $\rho=1$ and larger range using $\rho=2$. Both hyper-parameter choices produce qualitatively good generated samples, but we noticed that choosing $\rho=2$ yields better FID scores (if sampling from a standard multivariate normal at $t_{0}=T$) and requires smaller training times. Results presented in the main text reflect the larger noise option. For completeness, we also showcase FID results for the smaller noise range networks in Table 6.
@@
Finally, all training was performed on single GPUs. As in [28], we report training duration in Mimgs and note that time needed to achieve 200 Mimgs is approximately 2 weeks on NVIDIA RTX 3090 or 4090 GPUs using hyperparameters shown in Tables 4, 5 . This training time can be substantially reduced (to approximately 2 days for 200Mimgs) if using same multi-GPU training settings as in [28].
@@
Table 5: CIFAR-10 & AFHQv2 Variable Training Parameters (Per Schedule)
Hyperparameter Name Dataset PRP PRR to 2% PRR to 10% PRR to 20%
Training Duration (Mimgs) CIFAR-10 275 300 450 450
Training Duration (Mimgs) AFHQv2 275 300 350 350
Exponential Inflation Constant ($\rho$) Both 2 1 1 1
Total (Nominal) Dimensions Both 3072 3072 3072 3072
Dimensions Kept Both 3072 62 307 615
$g_{preserved\_dims}$ Both 1 2 2 2
$g_{compressed\_dims}$ Both 1 0.98 0.89 0.75
Table 6: FID Values for PR-Preserving Networks Trained with $\rho=1$
Dataset FID Mean ±$\pm 2\sigma$
CIFAR-10 $23.14\pm 0.10$
AFHQv2 $23.15\pm 0.06$
@@ B.5.1 Roundtrip Experiments
For image datasets (CIFAR-10 and AFHQv2), we simulated full round-trips: integrating the pfODEs (6) forward in time to map original images into latent space and then backwards in time to reconstruct original samples. We run these roundtrips for a set of 10K randomly sampled images, three times per each schedule (PRP, PRR to 2, 10, and 20%) and compute pixel mean squared error between original and reconstructed images, averaged across the 10K samples. Values reported in Table 1 represent mean ±$\pm 2$ standard deviations of pixel MSE between these three different random seeds per each condition. For pfODE integration, we used the discretization schedule and Heun solver detailed above (Appendix B.3.2), with $t_{max}=15.01$, $\epsilon_{s}=10^{-2}$, and $N=118$ for all conditions.
@@ B.5.2 FID Experiments
For image datasets, we also computed Frechet Inception Distance (FID) scores [46] across 3 independent sets of 50K random samples, per each schedule (PRP, PRR to 2, 10, 20%). Values reported in Table 1 represent mean ±$\pm 2$ standard deviations across these 3 sets of random samples per each condition. Here again, we used the discretization scheme and solver described in Appendix B.3.2 with $t_{max}=15.01$, $\epsilon_{s}=10^{-2}$, and $N=256$ across all conditions. We chose $N=256$ here (instead of 118) because this provided some reasonable trade-off between improving FID scores and reducing total compute time.
@@
To obtain our latent space random samples $\mathbf{x}(T)$ at time $t_{0}=T$ (i.e, at the start of generation) we sample from a diagonal multivariate normal with either 1) all diagonal elements being 1 (for PR-Preserving schedule) or 2) all elements corresponding to preserved dimensions being 1 and all elements corresponding to compressed dimensions being equal to the same small value for a given PR-Reducing schedule (see Table 7).
@@
Table 7: CIFAR-10 & AFHQv2 FID Experiments Latent Space Compressed Dimensions Variance
Schedule Latent Space Compressed Dimensions Variance
PRR to 2% $2.15\times 10^{-7}$
PRR to 10% $5.6\times 10^{-8}$
PRR to 20% $6.8\times 10^{-9}$
B.6 Details of Toy Alpha-Shape/Mesh Coverage Experiments
To assess numerical error incurred when integrating our proposed pfODEs, we performed coverage experiments using 3D meshes and 2D alpha-shapes [72, 73] in select toy datasets (i.e., 2D circles and 3D S-curve). If our numerical integration were perfect, points initially inside these sets should remain inside at the end of integration; failure to do so indicates miscalibration of the set’s coverage. For these experiments, we sampled 20K test points from a Gaussian latent space with appropriate diagonal covariance. For PR-Preserving schedules, this is simply a standard multivariate normal with either 2 or 3 dimensions. For PR-Reducing experiments, this diagonal covariance matrix contains 1’s for dimensions being preserved and a smaller value ($10^{-2}$ for Circles, $2.5\times 10^{-3}$ for S-curve) for dimensions being compressed.
Next, we sampled uniformly from the surfaces of balls centered at zero and with linearly spaced Mahalanobis radii ranging from 0.5 to 3.5 (200 pts per ball). We then fit either a 2D alpha-shape (2D Circles) or a mesh (3D SCurve) to each one of these sets of points. These points thus represent “boundaries” that we use to assess coverage prior to and after integrating our pfODEs. We define the initial coverage of the boundary to be the set of points (out of 20K test points) that lie inside the boundary. We then integrate the pfODE backward in time (the “generation” direction) for each sample and boundary point. At the end of integration, we again calculate the mesh or 2D alpha-shape and assess the number of samples inside, yielding our final coverage numbers.
Similarly, we take our samples and boundary points at the end of generation, simulate our pfODEs forwards (i.e., the “inflation” direction), and once again, use 2D alpha-shapes and meshes to assess coverages at the end of this round-trip procedure. As shown in Figure 4 (panels (B), (C)), we are able to preserve coverage up to some small, controllable amount of error for both schedules and datasets using this process.
When performing PR-Reducing simulations in these experiments, we chose $g_{preserved\_dims}=1.15$ (circles, s curve) and $g_{compressed\_dims}=0.85$ (circles) or $g_{compressed\_dims}=0.70$ (S-curve). This was necessary in practice to avoid numerical underflow errors that arose from fitting 2D alpha-shapes or meshes to highly compressed (approximately low-rank) Gaussian latent spaces.

### v3

@@ 7 Discussion
Here, we have proposed a new type of implicit probabilistic model based on the probability flow ODE (pfODE) in which it is possible to perform calibrated, identifiable Bayesian inference on a reduced-dimension latent space via sampling and integration. To do so, we have leveraged a correspondence between pfODEs and diffusion-based models by means of their associated Fokker-Planck equations, and we have demonstrated that such models continue to produce high-quality generated samples even when latent spaces are as little as 0.03% of the nominal data dimension. More importantly, the uniqueness and controllable error of the generative process make these models an attractive approach in cases where accurate uncertainty estimates are required.
Limitations: One limitation of our model is its reliance on the participation ratio (7) as a measure of dimensionality. Because PR relies only on second-order statistics and our proposals (9) are formulated in the data eigenbasis, our method tends to favor the top principal components of the data when reducing dimension. However, as noted above, this is not simply a truncation to the lowest principal components, since dimensions still mix via coupling to the score function in (6). Nonetheless, solutions to the condition (8) that preserve (or reduce) more complex dimensionality measures might lead to even stronger compressions for curved manifolds (Appendix C.2.2), and more sophisticated choices for noise and rescaling schedules in (6) might lead to compressions that do not simply remove information along fixed axes, more similar to [66]. That is, we believe much more interesting classes of flows are possible. A second limitation is that mentioned in Section 3.2 and in our experiments: our schedule requires training DBMs over much larger ranges of noise than are typically used, and this results in noticeable tradeoffs in compression performance as the inflation gap and number of preserved dimensions are varied.
Related work: This work draws on several related lines of research, including work on using DBMs as likelihood estimation machines [50, 67, 31], relations with normalizing flows and hierarchical VAEs [67, 33, 68], injective flow models [21, 22, 23, 24, 25], and generative flow networks [69]. By contrast, our focus is on the use of DBMs to learn score functions estimates for implicit probabilistic models, with the ultimate goal of performing accurate posterior inference. In this way, it is also closely related to work on denoising models [51, 52, 66, 70] that cast that process in terms of statistical inference and to models that use DBMs for de-blurring and in-painting [71, 72]. However, this work is distinct from several models that use reversal of deterministic transforms to train generative models [73, 74, 75, 76]. Whereas those models work by removing information from each sample $\mathbf{x}$, our proposal relies critically on adjusting the local density of samples with respect to one another, moving the marginal distribution toward a Gaussian.
@@
Acknowledgments and Disclosure of Funding
This work was supported by NIH grants F30MH129086 (DdA) and 1RF1DA056376 (JMP).
We also thank Eero Simoncelli for comments and discussion on an early version of this work.
@@ B.1.4 Re-writing loss with optimal preconditioning factors
As highlighted in main text, for constant end integration time $T$ and $\rho$, the final scale ratio between preserved and compressed dimensions is dictated by the quantity $g_{*}-g_{i}$, which we called the inflation gap (IG). Higher inflation gaps (IGs) lead to more stringent exponential shrinkage towards zero in compressed dimensions (Tables 6, 11) and worse off generative performance (Table 2).
In PR-Reducing experiments, we set $\rho=1$ and constructed $\mathbf{g}$ by making all elements of $\mathbf{g}$ corresponding to preserved dimensions equal to 2 (i.e., $g_{preserved}=g_{max}=2$) and all elements corresponding to compressed dimensions equal to $g_{compressed}=g_{min}=g_{preserved}-\text{IG}$ (Tables 5, 10). Of note, for PR-Preserving experiments, all elements of $\mathbf{g}$ are set to 1 (i.e., $\mathbf{g=1}$, $\text{IG}=0$) and we chose $\rho=2$, such that all dimensions are inflated to the same extent and we match exponential constant used for preserved dimensions in PR-Reducing experiments.
@@ B.4.1 Toy DataSets
Toy models were trained using a smaller convolutional UNet architecture (ToyConvUNet) and our proposed preconditioning factors (Appendix B.1). For all toy datasets, we trained networks both by using original images as inputs (i.e., “image space basis”) or by first transforming images to their PCA representation (i.e., “eigenbasis”). Networks trained using either base choice were able to produce qualitatively good generated samples, across all datasets. For all cases, we used a learning rate of $10^{-5}$, batch size of 8192, and exponential moving average half-life of $50\times 10^{4}$. For PR-Reducing schedules, we set $\rho=1$ and constructed $\mathbf{g}$ as described in Appendix B.2 (Table 5). The only exceptions were networks used on mesh and HMC toy experiments (Appendices C.2.1, B.6), where we used instead $g_{preserved}=1.15$ across all preserved dimensions (circles, S-curve) and $g_{compressed}=0.85$ (circles), or $g_{compressed}=0.70$ (S-curve) - Table 5, $2^{nd}$ and $6^{th}$ rows. This yields a softer effective compression (i.e., smaller IGs) and is needed to avoid numerical instability in these experiments.
As explained in Appendix B.1, to construct our $\mathbf{c}_{noise}$ preconditioning factor, we sampled $t\sim\mathcal{U}(t_{min},t_{max})$, with $t_{min}=10^{-7}$ across all simulations and $t_{max}$ equal to the values shown in Table 4. In the same table, we also show training duration (in $10^{6}$ images (Mimgs), as in [49]), along with both the total number of dimensions (in the original data) and the number of dimensions preserved (in latent space) for each dataset and schedule combination. In Table 6, we showcase latent space (i.e., end of “inflation”) compressed dimension variances achieved for the different toy PR-Reducing experiments as a function of inflation gap (IG). As expected, higher IGs lead to more stringent shrinkage of compressed dimensions in latent space.
@@
Table 5: $g_{i}$ Values for Preserved vs. Compressed Dimensions for Toy Experiments.
Dataset Schedule Dimensions Kept IG $g_{preserved}$ $g_{compressed}$
Circles PRR 1 2.0 2.0 0.0
Circles PRR 1 0.3 1.15 0.85
Sine PRR 1 2.0 2.0 0.0
Moons PRR 1 2.0 2.0 0.0
S Curve PRR 2 3.0 2.0 -1
S Curve PRR 2 0.45 1.15 0.70
Swirl PRR 2 3.0 2.0 -1
Table 6: Toy Experiments Compressed Dimension Variance by Inflation Gap (IG)
Dataset Schedule Dimensions Kept IG Compressed Dimension Variance
Circles PRR 1 2.0 $4\times 10^{-7}$
Circles PRR 1 0.3 $1\times 10^{-2}$
Sine PRR 1 2.0 $4\times 10^{-7}$
Moons PRR 1 2.0 $4\times 10^{-7}$
S Curve PRR 2 3.0 $2\times 10^{-12}$
S Curve PRR 2 0.45 $2.5\times 10^{-3}$
Swirl PRR 2 3.0 $2\times 10^{-12}$
@@ B.4.2 CIFAR-10 and AFHQv2 Datasets
For our image datasets (i.e., CIFAR-10 and AFHQv2), we utilized similar training hyperparameters to the ones proposed by [49] for the CIFAR-10 dataset, across all schedules explored (Table 7). Shown in Tables 8, 9 are our specific choices for the exponential inflation constant ($\rho$) and training duration (in $10^{6}$ images - Mimgs) for the two main sets of experiments performed on image datasets, namely (1) experiments with constant inflation gap (IG=1.02) and varying the number of preserved dimensions $d$ on both datasets (Table 8), and (2) experiments with fixed $d$ ($d=2$) and varying inflation gaps for the AFHQV2 dataset (Table 9). Here, training duration was determined for each schedule based on when computed Frechet Inception Distance (FID) scores [65] stopped improving. We also showcase in Table 10 the specific values used for elements of $\mathbf{g}$ corresponding to preserved vs. compressed dimensions at different inflation gaps.
@@
Finally, training was performed in a distributed fashion using either 8 or 4 GPUs per each experiment (NVIDIA GeForce GTX TITAN X, RTX 2080) in a compute cluster setting. Generation (FID) and round-trip (MSE) experiments were performed on single GPU (NVIDIA RTX 3090, 4090, A5000, A6000). We report training duration in Mimgs and note that time needed to achieve 200Mimgs is approximately 2 days on 8GPUs (4 days on 4 GPUs) using hyperparameters shown in Tables 7, 8, 9. This is in agreement with previous train times reported in [49] using an 8 GPU distributed training set up.
@@
Table 8: Training Duration (in Mimgs) and Exponential Inflation Constant ($\rho$) for Dimension Reducing Experiments Using 1.02 Inflation Gap (IG) and Dimension Preserving Experiments (IG = 0.0)
Dataset Total Dimensions Dimensions Kept IG Training Duration $\boldsymbol{\rho}$
CIFAR-10 3072 1 1.02 300 1
AFHQV2 3072 1 1.02 250 1
CIFAR-10 3072 2 1.02 300 1
AFHQV2 3072 2 1.02 250 1
CIFAR-10 3072 30 1.02 300 1
AFHQV2 3072 30 1.02 450 1
CIFAR-10 3072 40 1.02 300 1
CIFAR-10 3072 62 1.02 250 1
AFHQV2 3072 62 1.02 450 1
CIFAR-10 3072 307 1.02 300 1
AFHQV2 3072 307 1.02 300 1
CIFAR-10 3072 615 1.02 450 1
AFHQV2 3072 615 1.02 450 1
CIFAR-10 3072 1536 1.02 300 1
AFHQV2 3072 1536 1.02 250 1
CIFAR-10 3072 3041 1.02 300 1
AFHQV2 3072 3041 1.02 200 1
CIFAR-10 3072 3072 0.00 275 2
AFHQV2 3072 3072 0.00 275 2
Table 9: Training Duration (in Mimgs) and Exponential Inflation Constant ($\rho$) for AFHQv2 Experiments Using Variable Inflation Gaps (IGs)
Total Dimensions Dimensions Kept IG Training Duration $\boldsymbol{\rho}$
3072 2 1.10 200 1
3072 2 1.25 250 1
3072 2 1.35 250 1
3072 2 1.50 200 1
Table 10: $g_{i}$ Values for Preserved vs. Compressed Dimensions at Different Inflation Gaps (IGs)
Inflation Gap (IG) $g_{\text{preserved}}$ $g_{\text{compressed}}$
1.02 2.0 0.98
1.10 2.0 0.90
1.25 2.0 0.75
1.35 2.0 0.65
1.50 2.0 0.50
@@ B.5.1 Roundtrip Experiments
For image datasets (CIFAR-10 and AFHQv2), we simulated full round-trips: integrating the pfODEs (6) forward in time to map original images into latent space and then backwards in time to reconstruct original samples. We run these round-trips for a set of 10K randomly sampled images, three times per each schedule investigated and compute pixel mean squared error between original and reconstructed images, averaged across the 10K samples. Values reported in Tables 1, 2 represent mean ±$\pm 2$ standard deviations of pixel MSE between these three different random seeds per each condition. For pfODE integration, we used the discretization schedule and Heun solver detailed above (Appendix B.3.2), with $t_{max}=15.01$, $\epsilon_{s}=10^{-2}$, and $N=118$ for all conditions.
@@ B.5.2 FID Experiments
For image datasets, we also computed Frechet Inception Distance (FID) scores [65] across 3 independent sets of 50K random samples, per each schedule investigated. Values reported in Tables 1, 2 represent mean ±$\pm 2$ standard deviations across these 3 sets of random samples per each condition. Here again, we used the discretization scheme and solver described in Appendix B.3.2 with $t_{max}=15.01$, $\epsilon_{s}=10^{-2}$, and $N=256$ across all conditions. We chose $N=256$ here (instead of 118) because this provided some reasonable trade-off between improving FID scores and reducing total compute time.
@@
To obtain our latent space random samples $\mathbf{x}(T)$ at time $t_{0}=T$ (i.e, at the start of generation) we sample from a diagonal multivariate normal with either (1) all diagonal elements being 1 (for PR-Preserving schedule) or (2) all elements corresponding to preserved dimensions being 1 and all elements corresponding to compressed dimensions being equal to the same small value for a given inflation gap (see Table 11).
@@
Table 11: Latent Space Compressed Dimensions Variance per Inflation Gap (IG), Both Datasets
Inflation Gap (IG) Latent Space Compressed Dimensions Variance
1.02 $2.15\times 10^{-7}$
1.10 $6.00\times 10^{-8}$
1.25 $6.80\times 10^{-9}$
1.35 $1.50\times 10^{-9}$
1.50 $1.76\times 10^{-10}$
For our CIFAR-10 comparison experiments against existing injective flow models, we used the same implementations for M-Flow [21], Rectangular Flows [22], and Canonical Manifold Flows [23] as in [23]. When training the comparison injective flows, we used the same hyper-parameters proposed in Appendix G.1 of [23] for the CIFAR-10 dataset. The only difference here is that we trained models with latent dimensions equal to $d=[30,40,62]$. Finally, comparison FID scores reported in Table 3 represent best score out of 3 independently generated sets, each with 10K samples. For our comparison models, inflation gap was fixed to $\text{IG}=1.02$ while $d$ was varied between 30, 40, and 62 and we utilized the same training hyper-parameters reported in Tables 7, 8. FID scores were computed using the same discretization, solver, and general set up described above.
B.5.3 FID and Roundtrip Integration Experiments with Additional Initialization Seeds
In the FID and roundtrip pfODE integration experiments in Tables 1, 2 we showcased variation arising from using different seeds when constructing initial generation or roundtrip samples. Another relevant source of variability in FID and MSE scores reported arises from different parameter initializations when training our denoiser networks. To assess this source of variability, we trained networks using three additional initialization seeds for our worse performing schedule at constant inflation gap ($\text{IG}=1.02$, PR-Reducing to 307 dimensions), for both datasets (AFHQv2, CIFAR-10). For each such initialization seed, we conducted similar FID and roundtrip integration MSE experiments as detailed in Sections B.5.1, B.5.2.
Tables 12, 13 showcase results for these experiments on each individual seed tested (top 4 rows) and aggregated across all seeds (bottom row). For top 4 rows, values reported represent mean ±$\pm 2\sigma$ of scores computed across three different sets of generation/roundtrip seeds (as in Tables 1, 2). For bottom rows, values reported represent mean ±$\pm 2\sigma$ of scores computed across all initialization and generation/roundtrip seeds (i.e., these represent mean ±$\pm 2\sigma$ over all aggregated experiments conducted for the given schedule and dataset).
Table 12: AFHQv2, FID and Roundtrip Experiments with Additional Initialization Seeds
Seed FID MSE
1 $16.41\pm 0.10$ $3.24\pm 0.16$
2 $15.63\pm 0.17$ $3.20\pm 0.15$
3 $16.53\pm 0.11$ $3.06\pm 0.21$
4 $15.64\pm 0.10$ $3.33\pm 0.13$
all (aggregated) $16.05\pm 0.85$ $3.21\pm 0.26$
Table 13: CIFAR-10, FID and Roundtrip Experiments with Additional Initialization Seeds
Seed FID MSE
1 $26.32\pm 0.07$ $1.04\pm 0.09$
2 $27.43\pm 0.15$ $0.97\pm 0.08$
3 $27.80\pm 0.19$ $0.84\pm 0.04$
4 $28.07\pm 0.13$ $0.71\pm 0.02$
all (aggregated) $27.41\pm 1.34$ $0.89\pm 0.26$
B.5.4 Additional Figures for FID and Round-Trip MSE Experiments on Image Benchmark Datasets.
Below we showcase additional examples of generated images and results of roundtrip pfODE integration for different schedules explored in our experiments. More specifically, Figure 6 showcases results of FID and roundtrip integration experiments on CIFAR-10 at constant inflation gap ($\text{IG}=1.02$) and varying number of preserved dimensions. Figure 7 showcases results of FID and roundtrip integration experiments done on the AFHQv2 dataset, with varying inflation gaps $\text{IG}=[1.10,1.25,1.35,1.50]$ and constant number of preserved dimesions ($d=2$).
Figure 6: Generation and Round-Trip Experiments for CIFAR-10 at IG=1.02 and varying number of preserved dimensions. Layout and setup same as for Figure 5 - see Appendices B.5.2, B.5.1 for details.
Figure 7: Generation and Round-Trip Experiments for AFHQv2 dataset with dimension reduction to 2D (PRR to 2D) at different inflation gaps (IGs). Top row: Generated samples for each inflation gap (IG) flow schedule (1.10, 1.25, 1.35, and 1.50), all with $d=2$. Bottom row: Results of round-trip experiments for same schedules. Leftmost columns are original samples, middle columns are samples mapped to Gaussian latent spaces, and rightmost columns are recovered samples.
B.6 Details of Toy HMC Experiments
As highlighted in Section 5, we utilized Hamiltonian Monte Carlo (HMC) [1, 56, 57, 58] to assess if errors in our network score estimates could result in mis-calibrated posterior distributions. In these experiments, we worked with the toy 2D circles dataset (using both PR-Preserving and PR-Reducing schedules) and began by constructing our observed data samples $\mathbf{x_{obs}}$ as follows: First, we sampled a set of latent variables $\mathbf{z}$ from a 3-component Gaussian Mixture Model (GMM) $p(\mathbf{z})=\sum_{i=0}^{2}w_{i}\mathcal{N}(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})$ with known means ($\boldsymbol{\mu}$), diagonal covariances ($\boldsymbol{\Sigma}$), and weights ($\mathbf{w}$) (Table 14). Second, we integrated the sampled $\mathbf{z}$ points backwards in time (“generation”) using our proposed pfODEs with score estimates taken from trained networks to obtain “noise-free” observed data samples $\mathbf{x_{nl}}$. Finally, we added a small amount of isotropic Gaussian noise to these samples ($n\sim\mathcal{N}(0,\sigma^{2})$, $\sigma^{2}=10^{-2}$), to obtain our final observed data, $\mathbf{x_{obs}}$.
Table 14: Ground-Truth Means, Covariance Diagonals, and Weights for Gaussian Mixture Model (GMM) Components Used in Toy HMC Experiments
GMM Component Schedule Mean Covariance Diagonal Weight
$0^{th}$ PR-Preserving $[0.0,0.0]$ $[5.625\times 10^{-1},5.625\times 10^{-1}]$ 0.50
$0^{th}$ PR-Reducing $[0.0,0.0]$ $[5.625\times 10^{-1},5.625\times 10^{-3}]$ 0.50
$1^{st}$ PR-Preserving $[-5\times 10^{-2},0.0]$ $[10^{-2},1.0]$ 0.25
$1^{st}$ PR-Reducing $[-5\times 10^{-2},0.0]$ $[10^{-2},10^{-2}]$ 0.25
$2^{nd}$ PR-Preserving $[5\times 10^{-2},0.0]$ $[1.0,10^{-2}]$ 0.25
$2^{nd}$ PR-Reducing $[5\times 10^{-2},0.0]$ $[1.0,10^{-4}]$ 0.25
We then used these observations, $\mathbf{x_{obs}}$, along with the HMC implementation provided in the hamiltorch library [56], to jointly sample from the posterior over ($(\{\mathbf{z}_{j}\},\mathbf{w})$, assuming $\{\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}\}$ known.
For both PR-Preserving and PR-Reducing experiments, we generated 2000 samples ($\mathbf{x_{obs}}$). For sampling, we used $L=15$ steps per sampling trajectory, discarding the first 500 samples as “burn-in.” Step sizes were $10^{-2}$ for PR-Preserving and $10^{-3}$ for PR-Reducing schedules. Because sampling required integration over the full generative trajectory and was slow to mix, requiring roughly 40 minutes per sample, we initialized our $\mathbf{w}$ and $\mathbf{z}_{j}$ estimates to ground truth values. In other experiments, we verified that other initializations quickly converged to these values, but this procedure avoided numerical instabilities associated with integration of the generative pfODE during the burn-in phase. Finally, to reduce sample autocorrelation, we thinned the resulting chains by a factor of 5.
As mentioned above, this procedure required multiple neural function evaluations (NFEs) for pfODE integration per HMC integration step, producing very long sampling times. For instance, using the single-GPU setup of hamiltorch required $\simeq 2$ weeks to pass burn-in for our PR-preserving schedule and $\simeq 4$ weeks for our PR-preserving schedule. As a result, sample numbers were somewhat small ($N=1872$, PR-preserving; $N=1635$, PR-reducing), and thinned traceplots still exhibited some considerable correlation (Figure 8), underscoring the impracticality of using sampling-based inference in these models.
Figure 8: Traceplots (post-thinning) for 3 random chains for PR-Preserving and PR-Reducing schedules. A: Traceplots for 3 random PR-Preserving chains, after thinning by a factor of 5. “X axis” represents sample number and “Y axis” represents value of zeroth dimension of sample ($\mathbf{\hat{w}_{0}}$).B: Same set up, only for 3 random PR-Reducing chains. Note that there is still some considerable correlation in the samples, even after thinning. Additionally, mixing is not particularly good.